# Comprehensive Review of Deep Unfolding Techniques for Next-Generation Wireless Communication Systems

### v3

@@ III-D Discussions and Takeaways
Deep unfolding has proven to be a compelling solution for signal detection in complex and high-dimensional problems, such as those in massive MIMO, low-resolution ADC, and time-varying systems. By transforming iterative algorithms such as gradient descent variants, ISTA, ADMM, and AMP into structured networks, unfolding approaches bridge interpretability and performance. Table III summarizes the novelties, performance gains, advantages and disadvantages of the most popular deep unfolding models for signal detection.
@@ IV-A Massive MIMO
In massive MIMO systems, where the channel matrix is often sparse, the majorization-minimization (MM) approach is effective for addressing sparse channel estimation [102]. In [103], the authors address the challenge of downlink channel estimation in massive MIMO. A two-timescale joint uplink/downlink dictionary learning and channel estimation algorithm is proposed. The short-term subproblems focuses on estimating sparse channel vectors for each timeslot. The MM algorithm is unfolded, and the sparse channel vectors are treated as trainable parameters. In contrast, the long-term subproblem involves updating the dictionary to adapt to dynamic environments for which a constrained stochastic successive convex approximation (SSCA) framework is employed without deep unfolding. Simulations results in [103] demonstrate that the proposed approach reduces pilot overhead and outperforms existing methods in both static and dynamic environments, particularly in environments with mobile users.
The matching pursuit algorithm is a greedy iterative method often applied in sparse channel scenarios like massive MIMO. It identifies channel components (basis functions) by matching them to the residual, which is the difference between the observed signal and the signal generated by the current channel estimate. Yassine et al. [104] introduced a deep unfolded algorithm, called mpNet, based on the classical matching pursuit algorithm. This network estimates the multiuser massive MIMO channel in an unsupervised manner, initializing with a dictionary of imperfect steering vectors. The system operates in time division duplex mode to leverage channel reciprocity. The network dynamically adjusts the number of layers based on incoming data, ensuring efficient performance in realistic environments.
@@
Sparse Bayesian learning is another effective approach for channel recovery [106]. It estimates the non-zero elements of the channel matrix while utilizing prior knowledge of sparsity. Hu et al. [107] introduced a deep unfolding framework called deep deterministic policy gradient (DDPG) for MU-MIMO systems. DDPG, a variant of deep Q-networks, is designed to handle problems in continuous action spaces [108]. Unlike conventional deep unfolding NNs with fixed depth, the authors leveraged DDPG to dynamically adjust the network depth based on the complexity of the problem. This adaptive design enhances the efficiency and performance of sparse channel recovery in varying scenarios. The channel estimation problem is formulated using an off-grid sparse Bayesian learning model. The sparse Bayesian learning algorithm is unfolded with trainable parameters grid points and off-grid variables. Simulation results show that the proposed method significantly outperforms fixed-depth models, achieving a 30% reduction in the number of layers while improving normalized MSE (NMSE) performance.
@@
Compressive sensing aided deep unfolding is explored in [111], where the authors apply Onsager correction for AMP as explained in [112]. The architecture of the unfolded model, learned AMP (LAMP) [111] is similar to unfolded ISTA in [20]. A learned denoising-based AMP network (LDAMP) is proposed in [113] to learn beamspace channel estimates. The authors demonstrate that the iterative sparse signal recovery algorithm with denoising achieves robustness in channel estimation particularly in scenarios with large antenna arrays and limited RF chains. Yi et al. in [114] introduced an unfolded network called learned AMP with deep residual learning (LampResNet). In this model, the AMP algorithm is unfolded to estimate the channel, and deep residual learning is employed to mitigate channel noise. NMSE analysis shows that the proposed network in [114] outperforms LAMP [111] and achieves performance comparable to LDAMP [113] in terms of estimation accuracy. However, exploiting channel sparsity does not always yield optimal estimation performance. To address this, the authors of [115] investigated the sparsity of the mmWave channel in the beamspace. They formulate the estimation problem as a sparse signal recovery task and improve convergence speed by unfolding the iterative PGD algorithm. In this approach, the step-size of the descent step is treated as a trainable parameter. In [116], an iterative trimmed-ridge regression algorithm is unfolded to improve beamspace channel estimation. The proposed algorithm involves parallel computation to achieve negligible latency. These algorithms were applied to both narrowband and wideband massive MIMO systems, to ensure versatility in channel estimation.
@@ IV-C Modern Multicarrier Modulation
Yiyun et al. [117] proposed a CSI recovery network for FDD downlink massive MIMO-OFDM systems. ISTA is unfolded and the network consists of compression and reconstruction modules for CSI compression and feedback. Building on ISTA, Hu et al. [118] introduced an unfolded algorithm called the learnable optimization and regularization algorithm (LORA) to enhance the precision of CSI feedback. This approach also unfolds ISTA, and the regularization term is learned through end-to-end training. The model captures CSI characteristics and adapts to different levels of channel estimation error. Furthermore, to improve CSI feedback accuracy and efficiency, Yangyang et al. [119] proposed an unfolded ISTA network. Their study showed that the real and imaginary components of the channel differential terms, i.e, the difference between CSIs at adjacent time slots, exhibit similar correlation patterns in both the angular and delay domains. This property enables a shared model to process both components, thereby reducing network complexity and enhancing the efficiency of the CSI encoder. Each of the networks proposed in [117, 118, 119] enhance CSI feedback performance for FDD massive MIMO-OFDM system. In [33], Guo et al. proposed a two-stage low-rank CSI feedback scheme for a low-rank OFDM mmWave channel. In massive MIMO systems operating in FDD mode, CSI is estimated by unfolding the fast ISTA. Fast ISTA is an improved version of the classical ISTA due to the introduction of momentum term that accelerates the convergence rate. The network, FISTA-Net, includes trainable step-size and thresholds to achieve accurate reconstruction performance. In [120], the authors explore quantized compressed CSI at the user equipment before feedback for massive MIMO-OFDM system. They propose a deep unfolding-based bit-level CSI feedback network, TiLISTA, to mitigate the effects of accumulated quantization errors. Building upon ISTA, they used tied learned ISTA as discussed in [121]. Since the trainable parameters are shared across all layers, the number of learnable parameters are reduced, offering a faster and more accurate solution to the signal reduction problem. FISTA-Net [33] focuses on faster convergence while TiLISTA [120] targets robustness to quantization errors.
@@
In [122], an unfolded sparse Bayesian learning algorithm is derived to capture channel sparsity in a mmWave massive MIMO-OFDM system with hybrid beamforming (HBF). The proposed network achieved a lower NMSE of $0.033$ outperforming the classical sparse Bayesian learning algorithm which recorded an NMSE of $0.1203$. Jianqiao et al. [123] proposed an unfolded inverse-free variational Bayesian learning framework to address the high computational complexity of matrix inversion in mmWave MIMO-OFDM systems. The sparse channel recovery performance is enhanced by learning the a priori parameters of the sparsely represented dictionary for channel estimation. In [124], based on approximate Bayesian inference and an extension of expectation-consistent approximation, the authors proposed a generalized expectation consistent algorithm, an iterative message-passing approach for sparse channel estimation. In a lens-based beamspace massive MIMO-OFDM system, this approach compensates for large channel attenuation in the mmWave band. The authors developed a deep unfolded unsupervised learning network called the learned denoising-based generalized expectation consistent signal recovery network. The generalized expectation consistent signal recovery algorithm as discussed in [125] is unfolded. Both the networks in [123] and [124] provide accurate channel recovery estimates for mmWave channel with few RF chains. In [126], the authors leveraged angle-domain channel sparsity to reduce uplink pilot overhead during channel estimation for mmWave MIMO-OFDM systems. They propose an unfolded multiple-measurement-vectors-based AMP network. The phase shift network and channel estimator are jointly trained, thus integrating the channel sparsity for improved performance.
@@
TABLE IV: Summary of typical deep unfolding models for channel estimation and their advantages and disadvantages.
Unfolding Model Main Novelties and Performance Gains Advantages and Disadvantages
LDAMP [113] Denoising-based unfolded AMP model for sparse-beamspace channel estimation in mmWave massive MIMO. Outperforms classical denoising AMP method with small number of RF chains at the receiver. Advantages: Simple architecture and fast convergence (within $5$ layers), require a few RF chains. Disadvantages: Fixed parameters post-training restrict the model’s adaptability, performance may degrade in varying scenarios.
LampResNet [114] Combines unfolded AMP with deep residual learning to enhance robustness against noise. Achieves close performance to LDAMP [113] at high SNR regime ($25$ dB). Advantages: Resilient to noise, consistent NMSE improvement under SNR variations. Disadvantages: Rely on prior information of beamspace channel, not consistent with realistic channel models, and network may fail to converge.
FISTA-Net [33] Fast ISTA-based deep unfolding for low-rank CSI feedback in mmWave OFDM systems. Outperforms the classical FISTA with lower complexity. Advantages: Fast convergence ($20$ layers). Disadvantages: Requires retraining for major environment variations.
mpNet [104] Deep unfold model based on the matching pursuit method for multiuser massive MIMO systems. Outperforms classical LS estimator in terms of NMSE under varying SNR. Advantages: Avoids ground-truth channel data by using unsupervised online learning on noisy LS estimates. Disadvantages: Need prior knowledge of physical wave propagation model, as small uncertainty of system parameters lead to high estimation performance loss.
LORA [118] The deep unfolding model based on ISTA introduces a learnable regularization module to adaptively capture the CSI structure. Improved CSI reconstruction accuracy by $>2$ dB NMSE over TiLISTA [117]. Advantages: LORA exhibits strong denoising capability under channel estimation error. Disadvantages: The number of layers is critical: using fewer than four layers degrades reconstruction accuracy, while increasing the number beyond four leads to higher computational complexity with only marginal improvements in NMSE.
@@ IV-D IRS
Tsai et al. [131] proposed a two-stage learned AMP network with row compression to address the joint estimation problem of direct and cascaded channels in IRS-aided mmWave systems. By utilizing the deep unfolding technique, the network combines the advantages of compressive sensing and DL. This approach enables reliable channel estimation with reduced training, which is crucial for efficiently reconfiguring wireless environments under IRS reflection control.
@@ IV-E IoT devices
Joint activity detection and channel estimation are critical in grant-free random access for massive machine type communication (mMTC), where numerous IoT devices access the network sporadically without scheduling. To enhance detection under imperfect CSI, Qiang et al. [31] proposed AMP-Net, an unfolded AMP based network. Each AMP iteration alternates between MMSE-based denoising and residual updates to minimize the MSE. The network uses four trainable scalar parameters to estimate activity probability, channel variance, regulate the denoiser, and control the state evolution. AMP-Net achieved accurate detection of the active device by extracting the CSI from the estimated device state matrix with a $4$ dB gain over classical AMP at SNR$=15$ dB. The authors in [132] proposed a grant-free access protocol for mmWave cell-free massive MIMO using a hybrid structure based on learned vector AMP, reducing reliance on prior channel knowledge and improving convergence over classical AMP. In [133], Ma et al. introduced a one-phase non-coherent scheme for mMTC. The data bits are embedded in pilot sequences to enable joint activity and data detection without channel estimation. They used unfolded AMP to leverage pilot correlation. In [134], an unfolded AMP with backpropagation was applied to an asynchronous random access scheme with data length diversity, addressing activity detection, channel estimation, and data recovery in mMTC.
@@ IV-F Discussions and Takeaways
Channel estimation in modern wireless systems, especially in massive MIMO and mmWave settings, poses distinct challenges due to high dimensionality, feedback overhead, and sparse propagation characteristics. Deep unfolding addresses these issues by embedding structural priors into learning pipelines, as demonstrated by models like LDAMP [113] and LampResNet [114]. Table  IV presents notable unfolded networks for channel estimation with their performance gains, advantages, and disadvantages.
@@ V-A Massive MIMO
In [146], learnable parameters are used to approximate matrix multiplication and inversion in WMMSE update rules. However, the network fails to exploit the inherent graph structure of wireless networks. This limitation is addressed in [148]. An unfolded WMMSE algorithm for transceiver design in multicell multi-user MIMO (MU-MIMO) interference channels with local CSI is proposed in [148]. The authors use graph convolutional networks (a variant of GNN) to unfold WMMSE, enabling coordination in multicell MU-MIMO systems. Graph filters and convolutional networks approximate WMMSE-related variables, such as weighted beamformer matrices, by leveraging graph learning techniques. However, the approach in [148] lacks flexibility in learning the exact transformation of WMMSE variables. To overcome this, [149] introduces a hybrid unfolded WMMSE algorithm for transmit beamforming in MU-MIMO ad-hoc networks. This method leverages GNNs to learn functional transformations of WMMSE parameters. Compared to the classical WMMSE algorithm, the proposed network reduces the number of layers, leading to faster convergence of the optimization problem. In [150], the joint optimization of precoders for multicell MU-MIMO systems is studied. The WMMSE algorithm is unfolded to enable distributed and synchronized precoding across cooperative BSs. The methods in [146, 147] rely on centralized optimization, leading to high complexity and backhaul overhead. In contrast, due to the distributed precoding, the proposed network in [150] reduces computational cost while maintaining efficiency. An unfolded WMMSE-based NN for massive MU-MIMO under imperfect CSI is studied in [151]. The proposed network enhances performance as compared to [144], with low complexity by leveraging CSI statistics for near-optimal precoder solutions, thus accelerating convergence.
@@
Gradient-based methods have been successfully applied to beamforming optimization with reduced complexity. However, these approaches can be further improved using advanced optimization frameworks like PGD or projected gradient ascent (PGA). A robust frequency-selective precoding algorithm that maximizes the expected weighted sum rate while considering per-antenna power constraints is developed for wideband MIMO-OFDM system in [153]. This frequency-selective precoding algorithm is then unfolded into a deep network, with the aid of PGD. The trainable parameters include step sizes and projections in each PGD layer. The proposed network achieves near-optimal sum rate with lower computational complexity. Zhu et al. [154] proposed a beamforming learning architecture by unfolding a parallel gradient projection algorithm for MISO heterogeneous systems. The trainable set includes the step-size for the gradient ascent update and weighted sum rate coefficients used to construct the beamforming gradient direction. The proposed network generalizes well across different numbers of antennas, different numbers of BSs, and different network sizes especially without retraining. PGA is a variant of the gradient ascent algorithm. It updates the solution in the gradient direction to maximize the objective function and then projects it back onto the feasible set. Lavi et al. [155] unfolded the PGA algorithm to optimize hybrid precoding for multiuser downlink MIMO, assuming perfect CSI availability. To handle noisy CSI, they implemented the projected conceptual mirror prox algorithm [156, 157]. The step-sizes for updating the analog and digital precoder matrices are tuned using mini-batch SGD, a data-driven approach. The solution is computed within a fixed number of iterations, defined as network layers, ensuring a balance between interpretability and efficiency. In [158], Bilbao et al. proposed an efficient analog beamforming design to maximize the achievable rate of massive MIMO in-band full-duplex systems. The underlying beamforming problem is non-convex, primarily due to constant modulus constraints imposed by the use of analog phase shifters. To overcome this challenge, they developed an unfolded algorithm based on PGA combined with ZF projection onto convex sets. The convex feasible set arises due to the requirements of the ZF constraint and unit norm beamforming vectors. This approach leverages the structure of the zero-forcing conditions and the constant modulus constraints, achieving significantly faster convergence and improved performance compared to classical iterative methods. Besides PGA and PGD, CGD is also an iterative algorithm that falls under the gradient-based optimization approach. In [159], the CGD algorithm is unfolded for constant envelope precoding to reduce computational overhead. The Riemannian manifold optimization (MO)-based CGD is implemented across network layers. Simulations show that this approach effectively suppresses multi-user interference, improving overall performance.
@@
A fast and robust precoding network for MU-MIMO systems with imperfect CSI and per-antenna power constraints is proposed in [160]. The objective function is a geometric program that minimizes the MSE between the transmitted and recovered symbol vectors at the user. The iterative optimization framework is derived from the MSE uplink-downlink duality-based algorithm, which is unfolded to design a robust MSE duality network (RMSED-Net). RMSED-Net is robust to channel estimation errors and achieves sum rate close to the upper bound of its benchmark.
In [161], the interior point method is unfolded for precoding design. It leverages interference in multi-user MISO systems to minimize power while meeting signal-to-interference and noise ratio (SINR) constraints by training the proximal operators and Lagrangian multipliers. The authors show that the execution time is reduced by ${}50-70\%$ compared to the classical method. However, the network is limited to phase shift keying (PSK) modulation, requiring reformulation for higher order and QAM modulations.
A constructive interference-based symbol-level precoding for multi-user MISO downlink transmission is discussed in [162]. The proximal Jacobian ADMM is unfolded. The unfolded network reduces average transmit power and execution time by 98% compared to proximal Jacobian ADMM. The ADMM algorithm struggles with nonlinear discrete optimization due to biased alternating updates. To address this, iterative discrete estimation (IDE) is introduced in [163], ensuring unbiased updates and better convergence. Wang et al. in [163] proposed an algorithm for downlink massive MU-MIMO systems with finite-alphabet precoding to minimize inter-user interference. To reduce hardware costs and power consumption in multi-user massive MIMO, He et al. [164] developed an unfolded network using the IDE-based algorithm in [163]. The unfolded network with learnable step-size and damping factor achieves lower BER across a range of SNR.
In [165] a learning-based channel semantic acquisition and beamforming solution is proposed for cell-free massive MIMO. The beamformer is constructed by unfolding successive over-relaxation-based linear beamforming algorithm which reduces estimation errors in HBF. The proposed approach reduces computational complexity and enhances robustness to imperfect CSI.
TABLE V: Summary of typical deep unfolding models for precoder design and their advantages and disadvantages.
Unfolding Model Main Novelties and Performance Gains Advantages and Disadvantages
Unfolded WMMSE [166] Pioneering deep unfolding model based on the WMMSE algorithm for downlink precoding in MU-MIMO systems. Achieves $10\%$ improvement in sum rate metric compared with the conventional WMMSE at SNR $=45$ dB. Advantages: The network has a few trainable parameters, leading to fast convergence. Disadvantages: Performance loss with ill-conditioned channel matrices because the matrix inversion in WMMSE is replaced by a trainable approximation.
IAIDNN [146] Unfolded WMMSE model for MU-MIMO precoding. Converges after $2.94$ seconds compared to $291.01$ seconds required in classical WMMSE and reduces WMMSE computational cost by up to $30\%$. Advantages: IAIDNN generalizes to varying system sizes (numbers of users and antennas) with minor adjustments. Disadvantages: High training computational complexity due to the increased number of trainable parameters.
PGA-Net [155] Unfolds PGA with trainable step sizes for hybrid precoding. Proposed network achieves better sum-rate than classical PGA with fewer iterations. Advantages: Fast convergence and generalize well with varying channel size, effective for wideband channels. Disadvantages: Lacks inter-user interference cancellation and involves high gradient computation complexity.
ManNet [181] Deep unfolding models for HBF (with both fully- and sub-connected architectures) based on Riemannian MO method. Over $600$ times faster and $6$ times complexity reduction compared to MO-AltMin algorithm [182] ,has lower complexity than PGA-Net [155]. Advantages: Sparsely connected network architecture, and thus, has small number of training parameters. Disadvantages: The dynamic RF chain-antenna connections has not been incorporated into the unfolding models.
@@
Lin et al. [167] studied a LMMSE transceiver design for uplink massive MU-MIMO with one-bit ADCs. To minimize the system MSE by optimizing the precoder at each user equipment, the PGD algorithm is unfolded. The authors focused on training only the step-size in each PGD iteration, significantly reducing the training overhead. In [168], the authors combined deep generative models with unfolding techniques to achieve near-optimal HBF while minimizing feedback and computational complexity. The receiver estimates the channel using GAN, enabling efficient optimization of digital and analog precoders. The HBF problem is formulated as a bilevel optimization. The algorithm unfolds an iterative bi-level optimization problem into the layers of the CNN. There are two levels are upper and lower layer. In the upper layer digital precoder is optimized given a fixed analog precoder while in the lower layer the analog precoder is optimized given a fixed digital precoder. This alternating iterative steps are then unfolded. The trainable parameter set include the weights of the CNN that maps digital precoder to the analog precoder. The unfolded bilevel optimization algorithm achieved near optimal HBF with low feedback and computational cost.

### v4

@@ III-D Discussions and Takeaways
Deep unfolding has proven to be a compelling solution for signal detection in complex and high-dimensional problems, such as those in massive MIMO, low-resolution ADC, and time-varying systems. By transforming iterative algorithms such as gradient descent variants, ISTA, ADMM, and AMP into structured networks, unfolding approaches bridge interpretability and performance. In Table III, we summarize the computational complexity, number of trainable parameters and convergence speed of the most popular deep unfolding models against those of classical methods for signal detection.
@@
Although the aforementioned unfolded detectors achieved improved performance with reduced complexity under both perfect and imperfect CSI, their robustness over practical hardware and synchronization impairments have not been investigated. The performance analysis in this direction for different unfolded detectors are worth investigating in future works.
@@ IV-A Massive MIMO
In massive MIMO, where channels exhibit sparsity, majorization-minimization (MM) is effective for sparse channel estimation [122]. In [242], the authors propose a two-timescale joint uplink/downlink dictionary learning and channel estimation framework. The short-term subproblem estimates sparse channel vectors by unfolding the MM algorithm and treating these vectors as trainable parameters. The long-term subproblem updates the dictionary using a constrained stochastic successive convex approximation (SSCA) method without unfolding. Simulation results show reduced pilot overhead and superior performance in both static and dynamic scenarios, especially with mobile users.
The matching pursuit algorithm is a greedy method used for sparse channel estimation, iteratively identifying channel components by correlating the residual with dictionary elements. Yassine et al. [219] proposed mpNet, a deep unfolded matching pursuit based network for unsupervised multiuser massive MIMO channel estimation, initialized with a dictionary of imperfect steering vectors. The network dynamically adjusts its number of layers based on the received data, providing efficient estimation in realistic settings. Building upon the unfolded network: mpNet [219], Chatelier et al. [27] have considered hardware impairments in the shifting of the subcarrier frequency to achieve SISO-OFDM channel estimation.
@@
Sparse Bayesian learning is another effective approach for channel recovery [93]. It estimates the non-zero elements of the channel matrix while utilizing prior knowledge of sparsity. Hu et al. [85] introduced a deep unfolding framework called deep deterministic policy gradient (DDPG) for MU-MIMO systems. DDPG, a variant of deep Q-networks, is designed to handle problems in continuous action spaces [148]. Unlike conventional deep unfolding NNs with fixed depth, the authors leveraged DDPG to dynamically adjust the network depth based on the complexity of the problem. This adaptive design enhances the efficiency and performance of sparse channel recovery in varying scenarios. The channel estimation problem is formulated using an off-grid sparse Bayesian learning model. The sparse Bayesian learning algorithm is unfolded with trainable parameters grid points and off-grid variables. Simulation results show that the proposed method significantly outperforms fixed-depth models, achieving a 30% reduction in the number of layers while improving NMSE performance.
@@
Compressive sensing aided deep unfolding is explored in [22], where the authors apply Onsager correction for AMP as explained in [48]. The architecture of the unfolded model, learned AMP (LAMP) [22] is similar to unfolded ISTA in [61]. A learned denoising-based AMP network (LDAMP) is proposed in [72] to learn beamspace channel estimates. The authors demonstrate that the iterative sparse signal recovery algorithm with denoising achieves robustness in channel estimation particularly in scenarios with large antenna arrays and limited RF chains. Yi et al. in [206] introduced an unfolded network called learned AMP with deep residual learning (LampResNet). In this model, the AMP algorithm is unfolded to estimate the channel, and deep residual learning is employed to mitigate channel noise. NMSE analysis shows that the proposed network in [206] outperforms LAMP [22] and achieves performance comparable to LDAMP [72] in terms of estimation accuracy. However, exploiting channel sparsity does not always yield optimal estimation performance. To address this, the authors of [243] investigated the sparsity of the mmWave channel in the beamspace. They formulate the estimation problem as a sparse signal recovery task and improve convergence speed by unfolding the iterative PGD algorithm. In this approach, the step-size of the descent step is treated as a trainable parameter. In [210], an unfolded iterative trimmed ridge regression algorithm is proposed for beamspace channel estimation. The method supports parallel computation for near zero latency and is applicable to both narrowband and wideband massive MIMO systems, ensuring versatile channel estimation performance.
@@ IV-C Modern Multicarrier Modulation
Yiyun et al. [202] proposed a CSI recovery network for FDD downlink massive MIMO-OFDM by unfolding ISTA, with compression and reconstruction modules for CSI feedback. Hu et al. [86] improved it with learnable optimization and regularization algorithm (LORA), by learning the regularization term via end-to-end training to adapt to different channel estimation errors. In [237], Yangyang et al. exploited the correlation between real and imaginary parts of CSI differential terms, enabling a shared model to reduce complexity and improve encoder efficiency. Each of the networks proposed in [202, 86, 237] enhance CSI feedback performance for FDD massive MIMO-OFDM system. Guo et al. [63] proposed FISTA-Net, unfolding fast ISTA with momentum, trainable step-sizes, and thresholds for faster convergence and accurate low-rank OFDM mmWave CSI feedback. For quantized CSI feedback, the authors in [25] introduced TiLISTA, a deep unfolding bit-level feedback network using tied learned ISTA, reducing learnable parameters and enhancing robustness to quantization errors. Building upon ISTA, they used tied learned ISTA as discussed in [124]. Since the trainable parameters are shared across all layers, the number of learnable parameters are reduced, offering a faster and more accurate solution to the signal reduction problem. FISTA-Net [63] focuses on faster convergence while TiLISTA [25] targets robustness to quantization errors.
@@
In [54], an unfolded sparse Bayesian learning method is proposed for mmWave massive MIMO-OFDM with hybrid beamforming, effectively capturing channel sparsity and reducing NMSE from $0.1203$ to $0.033$. Jianqiao et al. [29] introduced an unfolded inverse-free variational Bayesian learning framework to avoid costly matrix inversions, improving sparse channel recovery by learning the prior parameters of the dictionary. In [70], a generalized expectation consistent algorithm is unfolded for sparse channel estimation in lens based beamspace massive MIMO-OFDM, mitigating severe mmWave attenuation. In [29] and [70] accurate channel recovery with few RF chains is achieved. Additionally, in [133] the author proposed an unfolded multiple measurement vector AMP network that exploits angle-domain sparsity to reduce uplink pilot overhead, jointly training the phase-shift network and channel estimator for improved estimation performance.
@@ IV-D IRS
Tsai et al. [190] proposed a two-stage learned AMP network with row compression for joint estimation of direct and cascaded channels in IRS-aided mmWave systems. Using deep unfolding, the method integrates compressive sensing with learning to achieve reliable channel estimation with reduced training overhead, facilitating efficient IRS reconfiguration.
@@ IV-E IoT devices
Joint activity detection and channel estimation are critical in grant-free random access for massive machine type communication (mMTC), where numerous IoT devices access the network sporadically without scheduling. To enhance detection under imperfect CSI, Qiang et al. [160] proposed AMP-Net, an unfolded AMP based network. Each AMP iteration alternates between MMSE-based denoising and residual updates to minimize the MSE. The network uses four trainable scalar parameters to estimate activity probability, channel variance, regulate the denoiser, and control the state evolution. AMP-Net achieved accurate detection of the active device by extracting the CSI from the estimated device state matrix with a $4$ dB gain over classical AMP at SNR=15=15 dB. The authors in [111] proposed a grant-free access protocol for mmWave cell-free massive MIMO using a hybrid structure based on learned vector AMP, reducing reliance on prior channel knowledge and improving convergence over classical AMP. In [134], Ma et al. introduced a one-phase non-coherent scheme for mMTC. The data bits are embedded in pilot sequences to enable joint activity and data detection without channel estimation. They used unfolded AMP to leverage pilot correlation. In [13], an unfolded AMP with backpropagation was applied to an asynchronous random access scheme with data length diversity, addressing activity detection, channel estimation, and data recovery in mMTC.
@@
TABLE IV: Comparison of Classical vs Unfolded methods for Channel Estimation.
$T$: Layer, $N_{f}$: No. of subcarriers, $N$: No. of devices, $M$: No. of BS antenna, $m$: Size of transmitted signal vector, $n$: Size of measurement vector, $L$: Pilot length, $G$: No. of discrete angular domain grid of channel clusters.
Unfolded Computational Complexity Number of Parameters (Unfolded) Convergence speed (Number of layers)
Model Classical Unfolded Simulation Setup Classical Unfolded
LISTA-Net [32] ISTA: $\mathcal{O}(nm)$ $\mathcal{O}(Tnm)$ $n\times n+2T$ MIMO, $n=250$, $m=500$ 2020 iterations $14$ Layers
ADMM-Net [44] ADMM: $\mathcal{O}(N^{2}M)$ $\mathcal{O}(LNM)$ $3T$ Massive MIMO, $L=90$, $M=4$, $N=300$ $40$ iterations 2020 Layers
ComNet [56] LMMSE: $1.6$ mega FLOPs $0.37$ mega FLOPs $1.2$ MBytes OFDM, $N_{f}=64$, $64$-QAM, WINNER II channel model. $200$ iterations 55 Layers
SBL-CE-Net [54] Sparse Bayesian learning: $0.6115$ tera FLOPs $0.0189$ tera FLOPs $TG^{2}N_{f}$ MIMO, $N_{f}=8$, $N_{t}=N_{r}=12$, $G=48$ $100$ iterations 33 Layers
@@ IV-F Discussions and Takeaways
Channel estimation in modern wireless systems, especially in massive MIMO and mmWave settings, poses distinct challenges due to high dimensionality, feedback overhead, and sparse propagation characteristics. Deep unfolding addresses these issues by embedding structural priors into learning pipelines, as demonstrated by models like LDAMP [72] and LampResNet [206]. Table IV summarizes the computational complexity, number of trainable parameters, and convergence speed (in terms of layers required for convergence) of key unfolded algorithms for channel estimation. Additionally, studies reporting complexity in FLOPs are included to highlight the implementation cost.
@@ V-A Massive MIMO
In [83], learnable parameters approximate matrix operations in WMMSE, but the method does not exploit the graph structure of wireless networks. To address this, in [166] the authors unfolded WMMSE using graph convolutional networks for multicell MU-MIMO, enabling coordination with only local CSI, though with limited flexibility in learning exact transformations. In [41], the authors improved this by proposing a hybrid unfolded WMMSE that uses GNNs to learn functional mappings, reducing the number of layers and accelerating convergence. Distributed precoding across cooperative BSs was achieved in [221], lowering complexity and backhaul load compared to centralized approaches [83, 156]. Finally, in [195] unfolded WMMSE is applied to massive MU-MIMO with imperfect CSI by leveraging CSI statistics, achieving faster convergence and near-optimal precoding relative to the classical WMMSE algorithm [170].
@@
Gradient based methods are widely used for beamforming optimization due to their reduced complexity, and further gains have been achieved using advanced frameworks such as PGD and projected gradient ascent (PGA). In [214], a frequency selective precoding algorithm for wideband MIMO-OFDM was unfolded using PGD with trainable step sizes and projection parameters, achieving near-optimal sum rate at low complexity. Zhu et al. [248] unfolded a parallel gradient projection method for MISO heterogeneous networks, where trainable step sizes and weighted sum-rate coefficients enable generalization across antenna and network configurations. PGA itself performs gradient ascent followed by projection onto the feasible set. Lavi et al. [107] applied unfolded PGA to hybrid precoding, with trainable step sizes for analog and digital updates; for noisy CSI, they used a projected mirror-prox scheme. Bilbao et al. [21] unfolded PGA with ZF projection for analog beamforming under constant-modulus constraints in full-duplex massive MIMO, exploiting convex ZF-feasible sets for faster convergence and improved performance. They have considered the CSI error to account for hardware impairment and non linearity in the channel. Beyond PGD and PGA, CGD has also been unfolded in [79]. The authors incorporated CGD with Riemannian manifold optimization for constant-envelope precoding, reducing complexity while effectively mitigating multi-user interference.
@@
In [213], a fast and robust precoding network is proposed for MU-MIMO with imperfect CSI and per-antenna power constraints. The MSE minimizing geometric program and its uplink–downlink duality-based iterative solution are unfolded to create RMSED-Net, which offers robustness to channel estimation errors and achieves sum-rate performance close to the benchmark upper bound.
In [144], an unfolded interior-point method is used for symbol-level precoding in multi-user MISO, where trainable proximal operators and Lagrange multipliers enable power minimization under SINR constraints with a 50–70% reduction in runtime. Its main limitation is applicability only to PSK modulation. In [216], a constructive-interference-based precoding scheme unfolds the proximal Jacobian ADMM, yielding 98% lower transmit power and significantly reduced execution time. Since ADMM suffers from biased updates in discrete optimization, iterative discrete estimation (IDE) for unbiased convergence in finite-alphabet precoding is introduced in [194]. He et al. [75] unfolded this IDE-based algorithm for massive MU-MIMO, with learnable step sizes and damping achieving improved BER across SNRs. Finally, in [57] the authors proposed a learning-based channel semantic acquisition and beamforming framework for cell-free massive MIMO, where a successive over-relaxation-based beamforming algorithm is unfolded to reduce estimation error, lower complexity, and enhance robustness under imperfect CSI.
@@
Lin et al. [116] unfolded the PGD algorithm for LMMSE transceiver design in uplink massive MU-MIMO with one-bit ADCs, training only the step-size per iteration to minimize system MSE and reduce training overhead. In [15], deep generative models are combined with unfolding to solve HBF as a bilevel optimization problem. The upper layer optimizes the digital precoder given a fixed analog precoder, and the lower layer optimizes the analog precoder given a fixed digital precoder. Trainable CNN weights map between the precoders. This unfolded bilevel approach achieves near-optimal HBF with low feedback and computational complexity.